The sum of three numbers is 119. The second number is twice as big as the first. The third number is twice as big as the second. What are the largest of the three numbers?
Answer by addingup(3677)

x+y+z= 119
y= 2x
z= 2y
----------------
y= 2x multiply both sides by 2:
2y= 4x Since we said that:
z= 2y and:
2y= 4x then:
z= 4x
Let's write our equation with one variable:
------------------------
x+2x+4x= 119 Add x on left:
7x= 119 Divide both sides by 7:
x= 17 is your first number
2(17)= 34 is your second number
4(17)= 68 is your third and Largest number
Proof:
17+34+68= 119
119= 119 We have the correct answer